OpenStudy (anonymous):

For each function, what is the output of the given input? f(x)=4x+7, find f(4)

OpenStudy (anonymous):

you have to 4 on all x. f(4)= 4*4+7 F(4)= 16+7 F(4)= 23 thats the answer

OpenStudy (anonymous):

That's what I thought when I put the 4 in where x is but when I used the function keys on my calculator it came up with 16x+28. So I've confused myself

OpenStudy (anonymous):

yea that has happened to me too

OpenStudy (anonymous):

do I have to get ride of the 4 in the f(4)?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

you are welcome